Circular No. 8238

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003jv, 2003jw, 2003jx, AND 2003jy Further to IAUC 8237, R. Covarrubias reports the discovery of four

additional supernovae by the ESSENCE project:

SN  2003 UT  α_{2000}  δ_{2000}  R  z Type Age
2003jv Oct. 24 23^{h}27^{m}58.^{s}22 − 8^{o}57′11′′.8 22.2 0.4 Ia  + 3
2003jw Oct. 28 2 31 06.84 − 8 45 36.5 22.6 0.3 Ia  − 2
2003jx Oct. 30 2 06 33.40 − 3 48 39.6 23.6 0.2 II
2003jy Oct. 30 2 10 53.98 − 4 25 49.8 22.7 0.3 Ia  − 6

Covarrubias adds that each of the 17 supernovae is within 1′′–2′′ of the center of its host galaxy. Corrigendum: On IAUC 8237, the declination for SN 2003jn should read −9^{o}02′15′′.6 (the sign was given incorrectly).

NOVAE IN M31 Further to IAUC 8226, M. Fiaschi, F. Di Mille, and D. Tiveron report

the discovery, on Hα-filtered images taken with the 0.41-m reflector, of another apparent nova in M31 located at α = 0^{h}42^{m}44^{s}.43, δ = +41^{o}15′52′′.3 (equinox 2000.0), which is 1′′.2 east and 16′′.2 south of the galaxy’s center. Due to the proximity to the core of M31, the photometry is very difficult, but images from Oct. 25.942 and 27.992 UT yield mag ∼ 16 for the new object; nothing was visible at this located on a similar image taken on Oct. 15 (limiting mag 18).

K. Hornoch reports additional R magnitudes for the first nova reported on IAUC 8222: Oct. 25.062 UT, 18.1; Nov. 2.878, 18.3.

COMET C/2003 T3 (TABUR) Improved parabolic orbital elements from MPEC 2003-V47:

T = 2004 Apr. 28.9316 TT  ω = 43.^{o}7939
Ω = 347.0786 }2000.0
q = 1.479945 AU  i = 50.4310

Approximate total visual magnitude estimates (in moonlight): Nov. 3.91 UT, 11.5 (W. Souza, Sao Paulo, Brazil, 0.20-m reflector); 4.01, 11.8 (A. Amorim, Florianopolis, Brazil, 0.14-m reflector).

2003 November 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green